OpenStudy (anonymous):

Write an equation of the line given the following conditions: // to x-3y=-12 & passes through the point (6,5)

OpenStudy (anonymous):

// parallel mean same slope so same m first put it in normal style: x+12=3y => y=x/3+4 so the slope must be 1/3 now we have an equation of the form: y=x/3+q fill now with x-y coordinates of the point: 5=6/3+q => q=3 !! finally you have y=x/3+3 !! tadaaaa XD

OpenStudy (anonymous):

I don't understand @ all

OpenStudy (anonymous):

so, parallel mean that the equation have the same slope right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you only need to know what is the other number, the q of the y=mx+q equation

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nice, you have the m and the point sonstitute them in the y=mx+q equation and resolve! You'll get the q

OpenStudy (anonymous):

*sostitute

OpenStudy (anonymous):

I understand now thanks

OpenStudy (anonymous):

;) you're welcome!